OpenStudy (anonymous):

Will Give reward Please HELP Marie has $830 in her bank account and withdraws $60 each month. Denise has $970 in her bank account and withdraws $80 each month. In how many months will Marie and Denise have equal amounts of money in their accounts? A. 4 B. 5 C. 7 D. 9

OpenStudy (anonymous):

answer is c

OpenStudy (texaschic101):

830 - 60m = 970 - 80m --- add 80m to both sides 830 - 60m + 80m = 970 --- subtract 830 from both sides -60m + 80m = 970 - 830 -- combine like terms 20m = 140 -- divide both sides by 20 m = 140/20 m = 7 They will both be equal in 7 months 830 - 60m = 970 - 80m 830 - 60(7) = 970 - 80(7) 830 - 420 = 970 - 560 410 = 410 and they will both have $ 410 in their account

OpenStudy (anonymous):

maria 830 -60 denis 970-80 1. 770-890 2. 710-810 3. 650-730 4.590-650 5.530-570 6 470-490 7. 410-410 8. 350-330 9. 290-250 from the startng total you make a chart and subtract their withdrawls from each side to give an answer